# On diffusion-based generative models and their error bounds: The log-concave case with full convergence estimates

### v4

@@ C.2 Proof of the Main Result for the General Case
$\begin{split}&W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})\\ &\leq(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{M})2\sqrt{\epsilon}\\ &\quad+\sqrt{2}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{M})e^{-2\widehat{L}_{\text{MO}}(T-\epsilon)-\epsilon}\\ &\quad+\sqrt{2\zeta^{-1}}e^{(1+\zeta-2\widehat{L}_{\text{MO}})(T-\epsilon)}\sqrt{\varepsilon_{\text{SN}}}\\ &\quad+\sqrt{2}e^{2(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))(T-\epsilon)}\gamma^{\alpha}\sqrt{T-\epsilon}\\ &\quad\ \times\Bigg{(}\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\qquad\quad+4\zeta^{-1}M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\quad\quad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\end{split}$
@@
$\begin{split}&\qquad\quad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\qquad\times[M\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg{)}^{1/2}.\end{split}$ (74)
@@
The bound for $W_{2}(\mathscr{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})$ in (74) can be made arbitrarily small by appropriately choosing parameters including $\epsilon,T,\varepsilon_{\text{SN}}$ and $\gamma$. More precisely, for any $\delta>0$, we first choose $0\leq\epsilon<\epsilon_{\delta}$ with $\epsilon_{\delta}$ given in Table 3 such that the first term on the right-hand side of (74) is
@@
Next, we choose $T>T_{\delta}$ with $T_{\delta}$ given in Table 3 such that the second term on the right-hand side of (74) is
@@
Next, we turn to the third term on the right-hand side of (74). We choose $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ with $\varepsilon_{\text{SN},\delta}$ given in Table 3 such that
@@
Finally, we choose $0<\gamma<\gamma_{\delta}$ with $\gamma_{\delta}$ given in Table 3 such that the fourth term on the right-hand side of (74) is
@@
Using (75), (76), (77) and (78), we obtain $W_{2}(\mathscr{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})<\delta$. ∎
C.3 Proof of the Preliminary Results for the General Case
We provide the proofs of Section 3.2 and Appendix C.1.
Proof of Remark 3.5.
Using Assumption 3, we have
@@ C.3 Proof of the Preliminary Results for the General Case
Integrating and taking expectation on both sides in (79), using Young’s inequality and Remark 3.5, we have
@@
Proof of Lemma C.2.
Using (11), Lemma C.1 and Remark 3.5, we have, for any $t\in[0,T-\epsilon]$ and $p\in[2,4]$,
@@
The proof is completed using Assumption 3. ∎
Proof of Lemma C.4.
@@
Using Assumption 3, we have
@@
Taking $v=\frac{\nabla_{x}s^{(k)}(t,\theta,x)}{|\nabla_{x}s^{(k)}(t,\theta,x)|}$ in (80), we have
@@
Appendix D Table of Constants
Table 3 displays full expressions for constants which appear in Theorem 3.10.
@@ Appendix D Table of Constants
$C_{\mathsf{EM},2}(T)$ $O(Me^{T^{2\alpha+1}}T^{2\alpha+1}\widetilde{\varepsilon}_{\text{AL}})$ $e^{T(4+8\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))}(\mathbb{E}[|\widehat{Y}_{0}^{\text{EM}}|^{2}]+16\mathsf{K}^{2}_{\text{Total}}T(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})(1+T^{2\alpha})+2MT)\end{aligned}$

### v5

@@ D.2 Proof of the Main Result for the General Case
$\begin{split}&W_{2}(\mathcal{L}(Y_{t_{K}}^{\text{aux}}),\mathcal{L}(Y_{K}^{\text{EM}}))\\ &\leq\sqrt{\mathbb{E}\left[|Y_{t_{K}}^{\text{aux}}-\widehat{Y}_{t_{K}}^{\text{EM}}|^{2}\right]}\\ &\leq\sqrt{2}e^{2(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))(T-\epsilon)}\gamma^{\alpha}\sqrt{T-\epsilon}\\ &\quad\times\Bigg{(}\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\qquad\quad+4\zeta^{-1}M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\qquad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\\ &\qquad\quad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\qquad\quad\times[M\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg{)}^{1/2}.\end{split}$ (78)
Final upper bound on $W_{2}(\mathcal{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}}).$
Substituting 61, 65, 68, and 78 into 59, we have
@@
$\begin{split}&W_{2}(\mathcal{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}})\\ &\leq(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{M})2\sqrt{\epsilon}\\ &\quad+\sqrt{2}(\sqrt{\mathbb{E}[|X_{0}|^{2}]}+\sqrt{M})e^{-2\widehat{L}_{\text{MO}}(T-\epsilon)-\epsilon}\\ &\quad+\sqrt{2\zeta^{-1}}e^{(1+\zeta-2\widehat{L}_{\text{MO}})(T-\epsilon)}\sqrt{\varepsilon_{\text{SN}}}\\ &\quad+\sqrt{2}e^{2(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))(T-\epsilon)}\gamma^{\alpha}\sqrt{T-\epsilon}\end{split}$
×$\begin{split}&\quad\ \times\Bigg{(}\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\qquad\quad+4\zeta^{-1}M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\quad\quad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\\ &\qquad\quad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\qquad\times[M\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg{)}^{1/2}.\end{split}$ (79)
@@
The bound for $W_{2}(\mathcal{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})$ in 79 can be made arbitrarily small by appropriately choosing parameters including $\epsilon,T,\varepsilon_{\text{SN}}$ and $\gamma$. More precisely, for any $\delta>0$, we first choose $0\leq\epsilon<\epsilon_{\delta}$ with $\epsilon_{\delta}$ given in Table 4 such that the first term on the right-hand side of 79 is
@@
Next, we choose $T>T_{\delta}$ with $T_{\delta}$ given in Table 4 such that the second term on the right-hand side of 79 is
@@
Next, we turn to the third term on the right-hand side of 79. We choose $0<\varepsilon_{\text{SN}}<\varepsilon_{\text{SN},\delta}$ with $\varepsilon_{\text{SN},\delta}$ given in Table 4 such that
@@
Finally, we choose $0<\gamma<\gamma_{\delta}$ with $\gamma_{\delta}$ given in Table 4 such that the fourth term on the right-hand side of 79 is
@@
Using 80, 81, 82 and 83, we obtain $W_{2}(\mathcal{L}(\widehat{Y}_{K}^{\text{EM}}),\pi_{\mathsf{D}})<\delta$. ∎
D.3 Proof of the Preliminary Results for the General Case
We provide the proofs of Section 3.2 and Appendix D.1.
Proof of Remark 5.
Using Assumption 3.b, we have
@@ D.3 Proof of the Preliminary Results for the General Case
Integrating and taking expectation on both sides in 84, using Young’s inequality and Remark 5, we have
@@
Proof of Lemma 21.
Using 11, Lemma 20 and Remark 5, we have, for any $t\in[0,T-\epsilon]$ and $p\in[2,4]$,
@@
The proof is completed using Assumption 3.b. ∎
Proof of Lemma 23.
@@
Using Assumption 3.b, we have
@@
Taking $v=\frac{\nabla_{x}s^{(k)}(t,\theta,x)}{|\nabla_{x}s^{(k)}(t,\theta,x)|}$ in 85, we have
@@
Appendix E Table of Constants
Table 4 displays full expressions for constants which appear in Theorem 10 and Remark 12.
Table 4: Explicit expressions for the constants in Theorem 10 and Remark 12
@@ Appendix E Table of Constants
$C_{\mathsf{EM},2}(T)$ $O(Me^{T^{2\alpha+1}}T^{2\alpha+1}\widetilde{\varepsilon}_{\text{AL}})$ $&e^{T(4+8\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))}\\ &\quad\times(\mathbb{E}[|\widehat{Y}_{0}^{\text{EM}}|^{2}]+16\mathsf{K}^{2}_{\text{Total}}T(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})(1+T^{2\alpha})+2MT)\end{aligned}$
@@
$\widetilde{C}_{4}(T,\epsilon)$ $O(\sqrt{M}e^{T^{2\alpha+1}}T^{2\alpha+1}\widetilde{\varepsilon}^{1/2}_{\text{AL}})$ $&e^{(1+(3/2)\zeta+2\mathsf{K}_{3}(1+2T^{\alpha}))(T-\epsilon)}\\ &\ \times(\zeta^{-1/2}(T-\epsilon)^{1/2}C^{1/2}_{\mathsf{EMose},2}+8^{1/2}\zeta^{-1/2}(T-\epsilon)^{1/2}\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}\\ &\qquad\ +2\zeta^{-1/2}\mathsf{K}_{3}(1+2T^{\alpha})(T-\epsilon)^{1/2}C_{\mathsf{EMose},2}^{1/2})\end{aligned}$